### v1

# Forests of Uncertaint(r)ees: Using tree-based ensembles to estimate probability distributions of future conflict

Abstract
Predictions of fatalities from violent conflict on the PRIO-GRID-month (pgm) level are characterized by high levels of uncertainty, limiting their usefulness in practical applications. We discuss the two main sources of uncertainty for this prediction task, the nature of violent conflict and data limitations, embedding this in the wider literature on uncertainty quantification in machine learning. We develop a strategy to quantify uncertainty in conflict forecasting, shifting from traditional point predictions to full predictive distributions. Our approach compares and combines multiple tree-based classifiers and distributional regressors in a custom auto-ML setup, estimating distributions for each pgm individually. We also test the integration of regional models in spatial ensembles as a potential avenue to reduce uncertainty. The models are able to consistently outperform a suite of benchmarks derived from conflict history in predictions up to one year in advance, with performance driven by regions where conflict was observed. With our evaluation, we emphasize the need to understand how a metric behaves for a given prediction problem, in our case characterized by extremely high zero-inflatedness. While not resulting in better predictions, the integration of smaller models does not decrease performance for this prediction task, opening avenues to integrate data sources with less spatial coverage in the future.
Keywords Conflict forecasting $\cdot$ Uncertainty quantification $\cdot$ Machine learning $\cdot$ Armed conflict
1 Introduction
With the advent of fine-grained conflict data and the increasing availability of computational resources over the last decade and a half, significant efforts have been made to design conflict forecasting systems to warn policy-makers of emerging dangers and enable preventative action. These systems generally provide estimates of the probability of conflict occurrence or point estimates of conflict-related fatalities (see Rød et al. (2024) for a comparison of existing systems). However, Hegre et al. (2025) highlight two key problems limiting the utility of both such point predictions: First, they only reflect the most likely scenario. Second, they do not provide any confidence estimates for the predicted value. Importantly, both issues are closely linked, as the latter becomes more important the less likely the underlying scenario is. The proposed solution is to estimate a predictive distribution, which allows for the quantification of uncertainties.
The sources of uncertainty in conflict prediction can be broken down into uncertainties related to the nature of conflict, and uncertainties related to the (lack of) data used in conflict prediction. The former includes its complexity and related challenges of translating explanation to prediction (Ward et al., 2010), its relative rarity (Mueller and Rauh, 2022), its changing nature (Bowlsby et al., 2020; Hegre et al., 2021), and resulting doubts regarding the systematicity and therefore predictability of conflict (Chadefaux, 2017). The latter includes various data-related challenges in the field, such as concerns on quality, availability, and low resolution of data on potential predictors (e.g. Cederman and Weidmann, 2017; Mueller and Rauh, 2018), and a number of biases in conflict event datasets, many related to its reliance on news media (Weidmann, 2015; Althaus et al., 2022; Bazzi et al., 2022; Öberg and Yilmaz, 2025; Raleigh et al., 2023). Integrating this into a much larger body of research on uncertainty quantification (UQ) in machine learning (ML), one can also view these issues in terms of aleatoric and epistemic uncertainty, with the former resulting from inherently stochastic processes and therefore irreducible, and the latter related to data selection and the modeling process (Hüllermeier and Waegeman, 2021; Gruber et al., 2025). While it can be debated how much of the apparently stochastic nature of conflict (Chadefaux, 2017) is in fact the product of data issues, from a statistical perspective the various methods to quantify uncertainty can be loosely summarized as estimating the predictive distribution for a given observation (Gruber et al., 2025). This aligns with the approach proposed by Hegre et al. (2025) for conflict forecasts.
We design and evaluate a ML approach including UQ on the PRIO-GRID-month level (pgm, Tollefsen et al., 2012), combining various tree-based models. We specifically incorporate two key sources of data-related uncertainty: First, we rely on algorithms which are able to estimate distributions natively and for each pgm individually, accounting for local differences in data generation processes and conflict mechanisms. Second, we employ probabilistic hurdle ensembles to account for systematic biases between the recording of events and the recording of fatalities (Hegre et al., 2022b; Lacina, 2006). We select combinations of either Random Forest (Breiman, 2001) or XGBoost (Chen and Guestrin, 2016) classifiers with Distributional Random Forest (Cevid et al., 2022), Quantile Regression Forest (Meinshausen, 2006), or NGBoost (Duan et al., 2019) regression models through a custom AutoML [^1] AutoML refers to the automation of some or all components of machine learning, such as model or hyperparameter selection. For more information, see e.g. Hutter et al. (2019). setup. Additionally, we explore an avenue to address data availability issues and regionally varying conflict mechanisms by testing the integration of multiple regional models with limited geographical scope in a spatial ensemble. As a result, data coverage requirements for individual models are reduced, thus allowing for the incorporation of region-specific datasets. This yields three model specifications: a “global”-only model covering the whole area of interest, a “local”-only model combining all regional models, and a global-local model based on the best-performing individual components of each.
Our evaluation shows that all three approaches generally outperform benchmarks across three distributional metrics with only very few exceptions. The global model and the global-local combined model score very similarly, while the local model performs only slightly worse. However, in absolute terms the differences in scores are marginal, prompting further investigation. Since the zero-inflatedness of our target greatly reduces the informative range of our scores, we design an experiment with simulated data to examine the impact of varying the accuracy of distributional predictions on the Continuous Ranked Probability Score (CRPS), our main metric. Our results show that the marginal differences in absolute scores observed are likely the consequence of sizable differences in prediction quality.
We complement the standard aggregate evaluation approach with a ranking-based approach. We group the predictions by matching pgms with countries, evaluate each country group individually, rank the model and benchmarks according to their performance, and compare the mean ranks across all countries for each year and metric. While this approach disadvantages our model compared to the benchmarks for countries without violence, since our distributional approach assigns a non-zero base probability for violence in any context, it outperforms the benchmarks in countries where violence occurs, which is arguably more relevant to users of early-warning systems.
2 Sources of Uncertainty in Conflict Prediction
Upon closer inspection, predicting violence with reasonable certainty is a daunting task. The many challenges can broadly be grouped into two categories: the nature of conflict and its determinants, and the characteristics, biases and availability of the data used. The uncertainty resulting from these challenges means that point predictions, i.e. predictions without uncertainty estimates, often struggle to beat even simple heuristics on common evaluation metrics (Vesco et al., 2022).
Decades of research into the causes of armed conflict has identified a number of contributing factors from opportunity and feasibility to motivation and grievances (e.g. Fearon and Laitin, 2003; Collier et al., 2008; Collier, 2004; Blattman and Miguel, 2010), with differences between individuals and social groups (Humphreys and Weinstein, 2008; Østby, 2008; Cederman et al., 2013). However, this does not automatically mean these factors can be used to make reliable predictions for the future (Ward et al., 2010; Chadefaux and Schincariol, 2025). While they are known to increase the risk of conflict, no combination of these factors can be considered a sufficient condition for conflict, with very similar contexts resulting in widely different outcomes in practice. The complex nature of conflict together with the unpredictability of actors involved has even led some to question to what extent conflict is even systematic enough to be predicted (Chadefaux, 2017). The challenge is exacerbated by the relative rarity of conflict, especially at the pgm-level [^2] See 3.1: From 1990 to 2017, UCDP records violence in less than 0.4% of all pgms., which means in practice there are very few examples of conflict that can be used to derive very complex patterns from (Mueller and Rauh, 2022; Hegre et al., 2025). Uncertainty is further increased by continuously evolving situations on the ground and larger geopolitical shifts, which can affect underlying risk patterns and limit the validity of historic insights or change the conditions that might lead to violence (Cederman and Weidmann, 2017; Chadefaux, 2017; Bowlsby et al., 2020; Hegre et al., 2021).
Second, the data used to capture both conflict and the various potential explanatory factors suffers from multiple shortcomings and biases. For many socio-economic and political risk factors of conflict, data varies widely in quality (Cederman and Weidmann, 2017; Chadefaux, 2017; Murphy et al., 2024), e.g. due to difficulties and expenses connected to measurement, or deliberate misrepresentation by governments. For many countries, subnational and subyearly data is also either not available or lags too far behind reality to be useful for forecasting tasks at the pgm-level. Additionally, data on structural risk factors often shows too little variance for meaningful predictions regarding the timing of future conflict (Mueller and Rauh, 2018; Chadefaux and Schincariol, 2025), while other data is not available at all or only in limited contexts. Consequently, the only fine-grained, high-variance features included often are conflict history features, which tend to have an outsized impact in terms of predictive power (Hegre et al., 2021, 2022a; Mueller and Rauh, 2022; Chadefaux and Schincariol, 2025). On the one hand, this corresponds to the well-documented issue of conflict recurrence (Collier et al., 2003), on the other hand conflict likely also proxies many unobserved factors that have already resulted in violence. Existing forecasts therefore do much better at discovering locations at risk than predicting when conflict is most likely to erupt (Mueller and Rauh, 2018; Bazzi et al., 2022).
However, the conflict event datasets that forecasts rely on for the prediction target and the most important feature set also suffer from several biases and shortcomings themselves. Mostly based on news reporting, both the accuracy of the information contained and the selection of events covered are potential sources of bias (Earl et al., 2004; Althaus et al., 2022). While not perfect, information on the events ultimately recorded has been evaluated as reasonably accurate, with errors regarding exact location and number of fatalities within acceptable boundaries for the pgm prediction problem (Weidmann, 2015). In contrast, it is often hard to determine what portion of violence is reported at all and to what extent it is biased exactly, given the difficulty of establishing an accurate ground truth in a conflict context (Price and Ball, 2015; Althaus et al., 2022). Selection biases generally depend on the widely varying judgments of newsworthiness and (partisan) preferences by a particular source (Davenport and Ball, 2002; Baum and Zhukov, 2015; Dietrich and Eck, 2020), while also impacted by logistical factors such as the possibility for information to reach reporters (Weidmann, 2016; Croicu and Kreutz, 2017). In sum, this likely results in considerable variance in data quality across different contexts.
These limitations add up to significant hidden uncertainty in point predictions, with forecasts for a given month and location almost guaranteed to be wrong to some degree.
2.1 Uncertainty Quantification
While UQ has received only limited attention in conflict prediction, a large body of research discusses UQ in machine learning settings. Uncertainty can be conceptually split into aleatoric uncertainty and epistemic uncertainty. Aleatoric uncertainty is caused by inherently stochastic processes and therefore irreducible, while epistemic uncertainty is the result of uncertainties related to the modeling process and can theoretically be minimized by additional data and better (fitted) models - although this may not always be possible in practice (Hüllermeier and Waegeman, 2021; Gruber et al., 2025). Methods to quantify this uncertainty can be loosely summarized as methods to estimate the predictive distribution, depending on the problem resulting either in probabilistic classification or prediction intervals (Cheng et al., 2023; Gawlikowski et al., 2023; Tyralis and Papacharalampous, 2024; Gruber et al., 2025), with many existing approaches rooted in Bayesian analysis (Berger, 1985; Lampinen and Vehtari, 2001).
Distinguishing between the types of uncertainty is often part of UQ approaches (Abdar et al., 2021; Hüllermeier and Waegeman, 2021; Gruber et al., 2025). However, Gruber et al. (2025) argue that data-related uncertainties blur the boundaries between aleatoric and epistemic uncertainty, with missing or biased data - a key challenge for conflict prediction - resulting in increased aleatoric uncertainty in practice, despite this being a theoretically solvable problem. In addition to what is discussed above, it could further be argued that the seemingly stochastic nature of conflict (Chadefaux, 2017) is simply the result of missing data, e.g. on the motivations of key actors, rather than an inherently stochastic process. Hence, instead of trying to disentangle these, focussing on the predictive distribution for UQ as proposed by the VIEWS team (Hegre et al., 2025) can be viewed as a sufficient first step in line with common practices for UQ in ML (Tyralis and Papacharalampous, 2024).
With these considerations in mind, we design an ML-based approach to conflict prediction incorporating UQ. In principle, ML is well-suited to capture the complexity of conflict through its ability to account for non-linear and interaction effects between multiple variables (Cederman and Weidmann, 2017). We use tree-based models, which often outperform the more popular deep learning approaches on tabular data (Grinsztajn et al., 2022). While deep learning approaches are in principle more capable in big data applications, their potential has so far not materialized in the field of conflict forecasts, likely due to the relative rarity of conflict occurrence. Our UQ approach specifically addresses two main data-related sources of uncertainty:
First, we exclusively use models with the ability to natively output uncertainty estimates. Importantly, given the potential of similar situations leading to widely different outcomes locally and over time, we focus on algorithms with the ability to predict distributions for each pgm individually rather than estimating uncertainty globally.
Second, we implement a hurdle approach to distinguish between zero and non-zero observations. This assumes differences in distributions on either side of a threshold, which is arguably the case in conflict data due to a combination of biases in the data-generation process as well as differences between the determinants of the intensity of conflict and the determinants of the occurrence of conflict (Lacina, 2006; Hegre et al., 2022b). The two-stage approach focusing first on the likely location of violence before predicting its intensity is also well suited to capture the spatial clustering and diffusion patterns of armed conflict (Buhaug and Gleditsch, 2008; Schutte and Weidmann, 2011; Racek et al., 2025).
Additionally, to work around some of the data availability challenges, we explore the selective combination of “local” and “global” models and its impact on model performance. While local predictions have been found to suffer from some of the same issues as the field of conflict forecasting as a whole (Bazzi et al., 2022), their integration into larger spatial ensembles is yet untested. Our contribution thus paves the way for the selective inclusion of locally available data sources into larger forecasting systems.
3 Methods
As the model was submitted to the VIEWS prediction challenge (Hegre et al., 2025), we generated predictions for each month in six yearly test windows (2018-2023) and (at the time) true future predictions for July 2024-June 2025.333The true future predictions are tracked and evaluated at https://viewsforecasting.org/research/prediction-challenge-2023/leaderboard/.
3.1 Data and Modeling Setup
With our main focus on modeling strategies, we rely on the data provided by the VIEWS (Violence & Impacts Early-Warning System) team in the context of the challenge and use all features provided [^4] These include information on conflict history, geography, natural resources, population size, climate and vegetation, and to some extent the economy. A codebook is available with the data provided in the context of the VIEWS challenge (Hegre et al., 2023). in our models. The data covers all PRIO-GRID cells in Africa and the Middle East, a subset with N=13110 cells of the global 0.5° x 0.5° grid (Tollefsen et al., 2012), and are available monthly starting in 1990. As outlined in the competition call, the target is the number of fatalities from state-based armed conflict events (Hegre et al., 2023), as recorded by the Uppsala Conflict Data Program (UCDP) (Sundberg and Melander, 2013; Davies et al., 2025). The target is highly zero-inflated, with less than 0.4% non-zero values in the average month in our training data.
To generate predictions for the whole next year from the available training data, we chose to train separate models for each of the timesteps to predict (t+3, … t+14) and combine the resulting outputs to a full year of predictions. The data cutoff for each test window is 3 months before the start, i.e. October 2017 for 2018 predictions. [^5] For each of the six yearly test windows, the training data is artificially limited to October of the previous year. This is done to simulate the availability gap present in the data for the true future predictions, where training data was available until April 2024 to predict for the timeframe July 2024 to June 2025. Therefore, predictions need to be generated for the timesteps t+3, …, t+14. The training data is further limited by the time period we want to predict into the future with this approach. For example, the model generating the predictions for December 2018 (t+14) based on data up to October 2017 (t) can only be trained on data up to August 2016 (t-14), as this is the last month where there is sufficient future information to label the target.
We designed a modular, model-agnostic modeling pipeline in Python, which performs the tuning, training and predicting automatically for the given prediction problem. We include multiple machine learning algorithms, selecting the algorithm which achieves the best performance during cross-validation for each timestep individually. This allows us to integrate and test different machine learning algorithms with minimal effort and means our approach can be easily reused for different prediction problems, with the code available as part of our replication material.
We perform hyperparameter tuning for each timestep once based on the data up to October 2017 (N=4,378,740), before the first test window. Our tuning procedure is based on time series cross-validation with a 5-year sliding window through the training data. We employ the hyperopt package (Bergstra et al., 2013), which implements a Bayesian search approach with the Tree-structured Parzen Estimator (TPE) algorithm (Bergstra et al., 2011), to identify the best hyperparameters for each prediction timestep model. The cross-validation and scoring within the tuning procedure is implemented in the scikit-learn framework (Pedregosa et al., 2011). After tuning, we fit the model on all available training data for a given test window before generating our predictions from the last available observations 3 months before the start of the prediction window.
Our tuning metric depends on the modeling step in the hurdle model, with further details provided below. For each chosen base metric, our tuning metric is defined as the mean performance across the test folds during cross-validation. As we observed a strong tendency to overfit for some parameter combinations, we combine this with a penalty term on deviations between mean train and test fold performance, in order to favor generalizability of the models. [^6] The formula we use is $M_{tune}=\overline{M}_{test}-0.5|\overline{M}_{train}-\overline{M}_{test}|$, where higher scores for all $M$ are better, and $\overline{M}$ represents the mean performance across all train respectively test folds in the cross-validation. The tuning metric $M_{tune}$ is then maximized by the TPE tuning algorithm.
3.2 Hurdle Approach
Our overarching modeling approach is a variation on the principle of hurdle models. Hurdle models (Mullahy, 1986) are a combination of two modeling steps: a first step consisting of a classifier determining whether the hurdle is reached, trained on all available training data, and a second step consisting of a regressor determining the predicted value, trained on only the subsection of the training data where the target has reached the hurdle. We perform tuning and prediction for each of these two steps separately to allow for custom combinations of global and local predictions (see below).
3.2.1 Classifier
Our classifiers are trained on a binary variable indicating whether or not any fatalities occurred in a given pgm. We include two different probabilistic classifiers in our modeling pipeline, Random Forests (Breiman, 2001) and eXtreme Gradient Boosting (Chen and Guestrin, 2016) models. [^7] We also tested logit models as a “simple” alternative approach, which performed worse by a factor of 2-3 on average.
  • Random Forests (RF) estimate probability by aggregating the results of a multitude of decision trees. Each tree in the forest is built from a different sample of data, using a technique called bootstrap aggregating, or bagging. Additionally, Random Forests employ random feature selection, where each split in a tree considers only a random subset of features.
  • eXtreme Gradient Boosting (XGB), employs a regularized learning objective that balances model complexity and predictive accuracy. The system utilizes gradient tree boosting, where the model is trained in an additive manner, incrementally improving the predictions by minimizing a loss function using second-order gradient statistics. Models are built sequentially, correcting the errors of previous models.
Our tuning metric of choice for classification is the average precision score, which summarizes the precision-recall-curve. This metric is well suited for zero-inflated classification tasks as it does not take into account whether zeroes are predicted correctly, which we argue is of little interest given the relative scarcity of violence (also see Saito and Rehmsmeier, 2015). Tuning performance for both model types is fairly similar, with RF performing slightly better in all global models, while XGB is favored for most local models.
3.2.2 Regressor
To estimate uncertainty around predictions, we rely on regressors designed to output distributions directly rather than trying to estimate distributions around point predictions ex post. We include three different tree-based distributional regressors in our modeling pipeline, Quantile Regression Forests (Meinshausen, 2006), Distributional Random Forests (Cevid et al., 2022) and Natural Gradient Boosting for probabilistic regression (Duan et al., 2019).
  • Quantile Regression Forests (QRF) extend the RF methodology to estimate conditional quantiles. Like RF, the QRF algorithm involves growing an ensemble of decision trees using a randomized node and split point selection process. Unlike traditional RFs, which only retain the mean response in each leaf, QRF retains all observed responses, enabling the estimation of the entire conditional distribution. The algorithm calculates the conditional quantile by averaging the weighted distribution of observed responses, with weights derived from the original RF methodology. We use evenly spaced quantile steps to generate the samples for our predictions with this algorithm.
  • Distributional Random Forests (DRF) are another extension of the traditional RF framework used to estimate the entire conditional distribution of univariate or multivariate responses. The methodology involves constructing trees that split data points based on a novel criterion derived from the Maximum Mean Discrepancy (MMD) statistic, which measures differences in distributions rather than just differences in means. This splitting criterion is applied recursively to ensure that the distributions in the resulting child nodes are as homogeneous as possible. Each tree in the forest is grown to optimize this distributional metric. The final forest model uses a weighted combination of trees to estimate the full conditional distribution of the response variables. This approach allows DRF to adaptively weight training data points based on their relevance to the prediction, providing a robust and flexible method for modeling complex dependencies.
  • Natural Gradient Boosting (NGB) for probabilistic regression extends gradient boosting to the estimation of probability distributions. This involves boosting the parameters of a specified parametric distribution using a natural gradient, which corrects the training dynamics for more stable and efficient learning. The algorithm integrates three modular components: a simple base learner, a parametric probability distribution, and a proper scoring rule. The natural gradient is employed to optimize the parameters of the conditional distribution, ensuring that the updates are invariant to reparameterization and efficiently exploit the curvature of the score in distributional space. We use decision trees as the base learner, log-normal probability distributions and the CRPS as the scoring rule.
Following the principle of hurdle models, we train our regression models only on pgms with non-zero targets while still generating predictions for all pgms. As our tuning metric, we use the competition’s main metric, the CRPS (Hegre et al., 2025). In line with the maximum number of samples allowed by the prediction challenge, we set our regression models to output 1000 samples of the predicted distribution. This results in a wider range of possible values, ensuring the inclusion of low-probability outcomes. NGB performed best during tuning, being chosen in 75% of global models and 80% of local models, with the other two regressors only chosen occasionally.
3.2.3 Quasi-Hurdle Ensemble
Hurdle model point predictions are usually generated via a simple multiplication of the output of the classifier and the predicted value of the regressor. Given that we work with N=1000 samples drawn from the predictive distribution instead, a multiplication of the classification probability with each of the samples would result in non-integer predictions [^8] The only exception here is NGB which does also predict zeroes. For consistency with the other algorithms we replace all zeros with ones in the NGB predictions., which is not in line with the nature of fatality counts. At the same time, our tree-based regressors trained only on non-zero targets never produce zero predictions and a multiplication would therefore likely overestimate the probability of violence occurring. While both issues could be partially addressed with rounding, we opt to instead interpret the classification probability as the percentage of the ensemble sample taken from the non-zero predictions via a random draw, with the remaining share of the 1000 samples filled with zero values. In testing, this also performed better than the multiplicative approach. [^9] We also tested selecting either all-zero samples or the full non-zero sample based on the predicted probability from the classifier and a threshold, which performed significantly worse.
3.3 Local Models
In addition to the “global” approach described above, we explore the integration of multiple “local” models encompassing smaller geographic regions. This has two potential benefits: first, it reduces the coverage requirements and thus increases the pool of available datasets for a prediction problem. Second, it accounts for potential biases across different regions, either due to differences in conflict patterns, or due to differences in data generation. Focussing on potential differences in conflict patterns, we create custom contiguous geographic clusters based on the spatial distribution of grid cells with any recorded fatalities in the training data (1990-2017) using the HDBSCAN clustering algorithm (Campello et al., 2013) [^10] We tested two additional versions of generating clusters: One with clusters created via an alternative clustering algorithm, DBSCAN, and a similar manual tuning of clustering parameters, and one with clusters corresponding to the United Nations Statistics Division sub-regions more aligned to the data coverage problem, with grid cells assigned to countries based on a majority rule. Both performed only slightly worse in testing. While we did not include them in our final model run and evaluation, corresponding models can still be produced with our replication code.. We manually tested parameter combinations for the clustering algorithm until inspections of the results yielded groupings, which plausibly corresponded to visually discernible patterns, resulting in eleven clusters. To ensure sufficient non-zero training data in each cluster for the hurdle regression models, we further reduce this down to six clusters by iterating over the clusters and combining smaller clusters with their nearest neighbors based on centroid distance of polygons drawn around each cluster, requiring a minimum of 1000 pgms with non-zero fatalities.
To subsequently assign any cells not containing conflict, we first draw new polygons around the combined grid cells of each cluster. Cells remaining outside these polygons were assigned to the cluster with the nearest boundary to the cell center. The procedure and resulting clusters are visualized in Figure 1. We subsequently train separate “local” models for each of the clusters following the same procedure as with the “global” models, described above. Each grid cell is therefore assigned not only to exactly one cluster but also to one corresponding local set of models. Combining the predictions from all local models yields predictions for the whole geographical area of interest.
3.3.1 Global-Local Ensemble
We create two spatial ensembles in addition to the original global hurdle model: a local-only hurdle model following the same approach, simply concatenating the predictions from all local models along their spatial IDs, and a global-local ensemble. The latter is constructed by selectively picking and choosing both classification and regression components from either the global or the local model, creating an ensemble of global and local model outputs. We select the components by comparing the combined performance of the hurdle ensemble for each of the four possible combinations of classification and regression predictions on a cluster-by-cluster basis, selecting the global-local combination for each cluster which performed best across the three years prior to a given prediction window. [^11] We also tested this using only one or two years of prior data, which resulted in worse performance of the combined prediction. This is likely connected to a fairly high volatility in performance across years, with prior performance not correlated enough to future performance. An increase to five years of prior data as the selection basis did not lead to meaningful improvements. Finally, the resulting predictions for each cluster are spatially concatenated analogous to the local-only model.
Figure 1: Visualization of the creation process of the clusters for local models. a) clusters created by manually tuned HDBSCAN algorithm with corresponding polygons; b) clusters after merging of smaller clusters with updated polygons; c) final clusters with grid-cells without violence assigned. Grid cells shown in a) and b) are those experiencing any violence in the training data (1990-2017). Black grid cells displayed in a) and b) are not assigned to any clusters by HDBSCAN initially.
4 Results
To illustrate what kind of intervals our model predicts, Figure 2 compares observed fatalities during the first test window with our predictions for four example grid cells representing typical prediction patterns: A high-intensity case (a) and a low-intensity case (b) we capture fairly well, and two “misses”: a “false negative” case (c) where our predictions showed only a minuscule chance for violence but in fact a significant number of fatalities occurred, and a “false positive” case (d) where our predictions saw a small but noticeable chance for violence usually seen in low intensity cases, but no violence occurred.
As is immediately apparent, our predictions are heavily right-skewed with long tails and a number of zeroes, indicating mostly low probabilities in the classification step. Only small prediction intervals for the highest-intensity predictions have lower bounds above zero (roughly 25% and smaller, Figure 2a). Notably, these characteristics are also found in the distribution of observed fatalities across pgms. The predictions capture general conflict intensity well in most cases, with the distributional approach treating the spikes often observed in real-world data simply as lower likelihood events, but within the range of possible values. Typical misses include “false negative” cases, where significant violence occurred despite our sample containing more than 99% zeroes, such as in Cameroon in the vicinity of the border to Nigeria (Figure 2c). However, as also illustrated, our models always assume a small base probability of violence with the samples never being all-zero, which is the result of our hurdle combination approach and a non-zero minimum probability from the classification step. Lastly, we see a number of “false positives” (Figure 2d), where a sizable share of our samples is non-zero but this predicted risk does not materialize. These predictions generally are part of a spatial cluster of higher-likelihood predictions with fatalities observed only in some of its grid cells. The distributional approach thus highlights elevated risk in cells near active conflict and predictions are similar to those where low-intensity violence did occur (Figure 2b), without being able to reliably identify exact locations.
Figure 2: Prediction intervals for example grid cells based on 2018 predictions from the global model, compared to the observed number of fatalities and a maximum a posteriori (MAP) estimate for the predictive distributions. MAP estimates are calculated as the maximum of the probability density function estimated via Gaussian kernel density estimation. 25%, 50%, 75%, 90%, and 95% prediction intervals are only drawn if upper boundaries are not zero. Figures 2b) and 2c) also include the extreme boundaries (min/max) of the corresponding samples.
4.1 Aggregate Performance
To evaluate the overall performance of our three ensemble models [^12] For simplicity, we subsequently use the term “model” in this context to refer to the respective specification used to generate our global-only, local-only, or global-local ensemble prediction., we compare them to several benchmarks provided by the VIEWS team across the six yearly test windows. Those are two naive benchmarks and three “conflictology” benchmarks based on medium- to long-term conflict history. The naive benchmarks are samples drawn from a Poisson distribution centered around the last observed values for each grid cell and predictions with only zero values. The “conflictology” benchmarks all treat historic fatality counts as draws from the predictive distribution to generate forecasts. The first benchmark (“conflictology”) uses fatality counts from a specific grid cell during the previous 12 months for the respective prediction window (12 draws). The second benchmark (“conflictology neighbors”) follows the same principle, but uses the combined conflict history of the grid cell and its immediate neighbors (108 draws). The third benchmark (“bootstrap 240”) draws 1000 random samples from the grid cell’s conflict history of the last 240 months. All also adhere to the two-month gap between training and test data. For instance, the “conflictology” samples for all months in 2018 are based on the observed fatalities from November 2016 to October 2017 (Hegre et al., 2025). [^13] Information in addition to Hegre et al. (2025) based on the source code at https://github.com/prio-data/prediction_competition_2023/blob/a45796ce8d1ffdd82e879e05c46d90c58b460a66/benchmark.py.
The evaluation results for our models and the benchmarks are reported in Table 1. We base our evaluation on the challenge’s main metric CRPS and the two additional metrics specified in the prediction challenge, the Ignorance Score (IGN) and the Mean Interval Score (MIS), all of which evaluate samples from a distribution (Hegre et al., 2025). For all three scores, lower values are better and perfect predictions have a score of zero. For potential comparisons with point prediction models, we also report Mean Squared Error (MSE) and Mean Absolute Error (MAE) values based on point predictions calculated via maximum a posteriori (MAP) estimation. We use Gaussian kernel density estimation (KDE) to estimate the probability density function (PDF) for each sample, through which we select the maximum.
At first glance, our three models are reliably able to beat the benchmarks, with very few exceptions. Our global model performs slightly better than the local model, while the global-local combination performs about the same as the global model. Depending on the metric, it even pulls ahead in two to three out of the six test windows. While this means we are unable to exploit systematic differences across contexts with our approach to improve predictive performance, this still opens avenues for the selective inclusion of data with limited geographic availability.
Model Year     CRPS IGN MIS MSE MAE
Global 2018     0.1300 0.0681 2.1646 62.396 0.1487
Local 2018     0.1376 0.0824 2.4546 59.707 0.1516
Global-local 2018     0.1324 0.0691 2.2266 61.682 0.1538
All-zero 2018     0.1444 0.0916 2.8883 65.815 0.1444
Poisson (last) 2018     0.3860 0.1177 7.1488 169.45 0.4020
Conflictology 2018     0.1919 0.8589 2.8345 86.968 0.2232
Conf. neighbors 2018     0.1473 0.1770 3.0622 64.951 0.1583
Bootstrap 240 2018     0.1443 0.0925 2.8883 65.815 0.1444
Global 2019     0.1010 0.0664 1.5599 16.665 0.1153
Local 2019     0.1040 0.0805 1.8239 16.515 0.1156
Global-local 2019     0.1011 0.0668 1.6023 16.603 0.1173
All-zero 2019     0.1154 0.0944 2.3089 17.239 0.1154
Poisson (last) 2019     0.1442 0.1050 2.6166 18.224 0.1515
Conflictology 2019     0.1184 0.8561 1.8887 17.444 0.1275
Conf. neighbors 2019     0.1068 0.1755 1.8786 16.717 0.1204
Bootstrap 240 2019     0.1154 0.0951 2.3089 17.239 0.1155
Global 2020     0.1180 0.0743 1.9011 16.812 0.1311
Local 2020     0.1221 0.0872 2.2054 16.708 0.1338
Global-local 2020     0.1181 0.0742 1.9001 16.796 0.1315
All-zero 2020     0.1319 0.1077 2.6374 17.218 0.1319
Poisson (last) 2020     0.1646 0.1163 2.9928 18.747 0.1725
Conflictology 2020     0.1275 0.8599 2.0731 16.547 0.1365
Conf. neighbors 2020     0.1230 0.1818 2.1152 16.932 0.1334
Bootstrap 240 2020     0.1317 0.1072 2.6374 17.218 0.1319
Global 2021     0.9246 0.0843 17.9788 81844.5 0.9383
Local 2021     0.9293 0.0964 18.3496 81844.5 0.9405
Global-local 2021     0.9243 0.0843 17.9785 81844.2 0.9377
All-zero 2021     0.9398 0.1188 18.7961 81844.9 0.9398
Poisson (last) 2021     0.9703 0.1286 19.0801 81843.2 0.9793
Conflictology 2021     0.9302 0.8648 17.8700 81843.0 0.9426
Conf. neighbors 2021     0.9279 0.1893 18.1056 81844.5 0.9414
Bootstrap 240 2021     0.9396 0.1175 18.7961 81844.9 0.9398
Global 2022     1.1274 0.0834 22.2467 98555.7 1.1396
Local 2022     1.1289 0.0951 22.4550 98555.4 1.1402
Global-local 2022     1.1263 0.0832 22.2353 98553.7 1.1403
All-zero 2022     1.1375 0.1199 22.7494 98560.0 1.1375
Poisson (last) 2022     1.4565 0.1453 28.5266 98770.9 1.4734
Conflictology 2022     1.1419 0.8669 22.2770 98504.3 1.1532
Conf. neighbors 2022     1.1311 0.1896 22.4754 98547.9 1.1442
Bootstrap 240 2022     1.1373 0.1181 22.7494 98560.0 1.1375
Global 2023     0.2147 0.0863 3.9807 163.26 0.2275
Local 2023     0.2207 0.0988 4.2879 163.17 0.2565
Global-local 2023     0.2175 0.0867 4.1036 163.52 0.2444
All-zero 2023     0.2236 0.1210 4.4723 163.33 0.2236
Poisson (last) 2023     9.7500 0.1507 193.97 1134237 9.7807
Conflictology 2023     0.5237 0.8686 13.218 369.37 0.3576
Conf. neighbors 2023     0.2499 0.1922 4.0334 163.39 0.2338
Bootstrap 240 2023     0.2234 0.1196 4.4723 163.33 0.2236
Table 1: Overview of model and benchmark metrics. Best results for each year are marked in bold. Lower scores signify better performance. Note that the CRPS is always equal to the MAE in the case of all-zero predictions.
4.2 Uncertainty in Model Evaluation
At second glance, differences in performance between both our own models and our models’ improvements over most benchmarks are miniscule, and could conceivably be caused e.g. by random components in the modeling process, warranting a closer look at the scoring functions and their properties. First, it should be noted that all three scores cannot be compared across different data and thus years. Yearly scores are calculated as the mean of the score for all $N$ pgms (Hegre et al., 2025) [^14] See https://github.com/prio-data/prediction_competition_2023 for the challenge-specific implementations of the MIS and IGN scores., e.g.
$\overline{CRPS}=\frac{1}{N}\sum_{i=1}^{N}CRPS(F_{i},y_{i})$ (1)
with
$CRPS(F_{i},y_{i})=\int_{-\infty}^{\infty}\left(F_{i}(x)-\mathbbm{1}(x-y_{i})\right)^{2}\,dx\quad\text{where}\quad\mathbbm{1}(z)=\begin{cases}1&\text{if }z\geq 0\\ 0&\text{otherwise}\end{cases}$ (2)
for the PDF $F_{i}(x)$ representing the prediction distribution and the corresponding observation $y_{i}$.
Both $CRPS_{i}$ and $MIS_{i}$ for all-zero predictions and no observed violence are 0, while $IGN_{i}$ is a fixed constant. Assuming most zeroes are predicted correctly, the highly zero-inflated nature of our prediction task means the informative range for each score becomes very small, as only a tiny percentage of predictions actually impact the sum, while N remains constant. This means the scores are not independent of the underlying data for a given application. For illustration, CRPS and MIS scores are almost perfectly correlated with the number fatalities across the different years, while IGN scores are strongly correlated with the number of non-zero pgms for most models and benchmarks (Table 2).
    CRPS     MIS     IGN
Model     Fatalities Non-zero pgms     Fatalities Non-zero pgms     Fatalities Non-zero pgms
Global     0.99998 0.66297    0.99990 0.66577    0.64328 0.98190
All-zero     1.00000 0.66025    1.00000 0.66025    0.66025 1.00000
Poisson (last)     -0.10729 0.54045    -0.10759 0.54067    0.52381 0.84532
Conflictology     0.96532 0.78414    0.91448 0.84714    0.60015 0.92618
Conf. neighbors     0.99957 0.67415    0.99951 0.64671    0.62442 0.97840
Bootstrap 240     1.00000 0.66016    1.00000 0.66025    0.65347 0.99991
Table 2: Correlation between evaluation scores and number of fatalities / number of non-zero pgms across the six test windows for global model and benchmarks. Note that the perfect correlation for all-zero samples with fatalities for CRPS and MIS, and with non-zero pgms for IGN is expected behavior. The perfect correlation of the Bootstrap 240 benchmarks with fatalities for the CRPS is the result of rounding, while for the MIS it is the result of the 90% prediction interval used (Hegre et al., 2025) in combination with the composition of the individual samples. As the samples never contain more than 1.9% non-zero draws for this benchmark in a given test window, the intervals only contain all-zero predictions.
Figure 3: Value ranges of CRPS scores for one year of simulated actuals and predictions, with varying noise and “accuracy”. Actuals are simulated with zero-inflatedness matching the training data, and values drawn from an estimated PDF based on non-zero observations < 1000 fatalities. Predictions are drawn from a poisson distribution with mean and variance equal to the corresponding actual. “Accuracy” ($\alpha$) less than 1 means a share of non-zero samples was replaced with all-zero predictions. Noise ($n$) means actuals were shifted by $50n\varepsilon$ [^15] While the cutoff point for the noise multiplier is chosen arbitrarily, the range should be sufficient to produce fairly strong effects given the values seen in our predictions - based on Figure 1 and a median number of fatalities of 5 in the training data. Increasing the multiplier much beyond $50\varepsilon$ continues to decrease the size of the CRPS change when varying the “accuracy”.before creating the samples, with $\varepsilon$ randomly drawn from a uniform distribution between -1 and 1.
Given this, it is essential to understand the informative range of the score to properly evaluate our models. We use simulated data to experimentally explore the impact of prediction accuracy for this degree of zero-inflation. For this, we focus on the CRPS as the main metric of the challenge. For our simulated observations $z_{i}$, we create a sample matching the size of one test window (N=157.320) and the share of non-zero pgms from the dataset (0.005, N=787). We draw simulated non-zero values $z_{i}^{+}$ based on a PDF estimated via Gaussian KDE from all observed values $x_{i}$ in the training data, where $0<x_{i}<1000$, thus controlling for extreme outliers, multiplying all resulting negative values with -1 to ensure we have valid fatality numbers. Next, simulated predictions are created based on these simulated observations. “Perfect” predictive distributions are created by drawing samples with size N=1000 from a Poisson distribution, setting the expectation $\lambda$ to the simulated actual value, i.e. $\lambda_{i}=z_{i}$ (see Hegre et al., 2023), thus obtaining a predictive distribution for each simulated observation $z_{i}$. Finally, we introduce errors to our predictions in two ways: First, we vary the “accuracy” $\alpha$ by replacing a share of the predictions for the simulated non-zero observations $z_{i}^{+}$ with all-zero predictions, i.e. for $\alpha=0.8$, 20% (157) of the 787 predictions for simulated non-zero observations are replaced with all-zero predictions. Second, we add different levels of random noise to all simulated non-zero observations before drawing the samples for our predictive distributions, i.e. $z_{i}^{+\prime}=z_{i}^{+}+\varepsilon$, randomly shifting the center of our predictive distributions away from the actual value. In both cases we still evaluate against the original $z_{i}$. [^16] We replace any values ¡ 0 with 0 before drawing the samples, to ensure the prediction samples only contain non-negative values.
Figure 15 shows the CRPS value ranges based on the simulation across varying “accuracies” and noise. CRPS values range from 0.004 to .101 going from “perfect” predictions to 0.1 “accuracy”, dropping on average .01 per 10% drop in “accuracy”. The more noise is added to the data, the smaller the change in CRPS becomes when varying the “accuracy”, with the changes more than half for the maximum level of noise in our simulation. Given that our best models are on average 0.007 points better than the best benchmark, this makes us confident that this means a real improvement in performance over the benchmarks rather than random variation.
Figure 4: Comparison of model-wide scores with scores based on grid cells grouped by individual countries for the three distributional metrics: CRPS (a), MIS (b), and IGN (c).
Finally, having confirmed that the small differences in our evaluation metrics matter, we want to understand where these differences come from. To do so, we match grid cells to countries based on a majority rule and evaluate our global model for each country individually. Country borders for assignment are taken from geoBoundaries (Runfola et al., 2020). Since this approach means strongly varying levels of zero-inflatedness across the countries, the corresponding scores also vary strongly and can no longer be directly compared (Figure 4). Instead, we rank the performance of all models and benchmarks for each country and calculate the average ranks across all scenarios (Table 3). Based on mean rank across all countries and years, our model comes in only 5th on the CRPS, 4th on the MIS, and 3rd on the IGN. This can be explained by roughly half of all countries never experiencing any violence, for which our model gets heavily penalized in a rank-based evaluation due to its base assumption of a small but non-zero chance of violence anywhere. [^17] For both the CRPS and the IGN, any non-zero values in our samples mean that predictions will be scored marginally worse. Since the benchmarks generally yield all-zero samples in countries without prior violence, this means our model will be ranked 6th in such cases, while the benchmarks all receive a joint 1st place. To illustrate: Out of 80 countries covered by our model, 32 never see any fatalities across all yearly test windows, with a minimum of 39 no-violence countries in any individual window. However, if we focus only on countries for which any violence at all was observed in a given year - what we argue is where the performance of our model matters most, our model ranks highest on average compared to all benchmarks and for all three distributional metrics.
    CRPS     MIS     IGN
Model     Mean rank Mean rank non-zero     Mean rank Mean rank non-zero     Mean rank Mean rank non-zero
Global     3.70 2.43    1.94 2.58    2.33 1.55
All-zero     2.08 3.24    1.81 2.68    1.82 2.70
Poisson (last)     2.60 4.24    2.24 3.51    2.12 3.28
Conflictology     2.38 3.55    2.07 2.91    5.88 5.75
Conf. neighbors     2.32 2.83    2.12 2.84    4.56 4.09
Bootstrap 240     4.85 3.91    1.81 2.68    3.55 3.14
Table 3: Rank-based evaluation of global model in comparison to benchmarks. Ranks were calculated for all countries based on the underlying grid cells per test window and averaged across countries and test windows. Grid cells were assigned based on the majority of the area within a grid cell covered. Mean rank includes all countries, while mean rank non-zero only includes countries with any fatalities recorded in a given test window. The highest non-zero mean ranks are marked bold.
5 Discussion
We demonstrate that it is possible to produce high-resolution conflict predictions with uncertainty estimates using a modeling architecture built from relatively simple components, which are able to beat a range of heuristic conflict benchmarks. We further find that these performance gains arise from cases where violence occurs, and thus in situations in which prediction tools are also most likely valuable to practitioners. With our evaluation, we highlight the need to not simply pursue marginal improvements in evaluation metrics, but to understand how they behave in the context of the prediction problem. While this is a useful first step towards actionable uncertainty estimates, the prediction intervals we produce are still fairly wide, with further research needed both to sharpen predictions and to reduce the sources of uncertainty.
Building on our approach, data-related avenues for improvement naturally include the use of additional data sources, potentially in combination with multi-level modeling to combine data at different resolutions, while also accounting for the distinction between macro- and micro processes of conflict (Balcells and Justino, 2014; Fritz et al., 2024). As our results show, it is also possible to combine multiple regional models without significantly losing performance. This opens up avenues to integrate data sources previously discarded due to coverage issues, e.g. collected by regional organizations such as ECOWAS. Additionally, data cleaning techniques such as outlier correction, applied both to the target and to the features, could reduce data-driven uncertainties. In addition to general improvements through better-suited methods, approaches such as selective classification (El-Yaniv and Wiener, 2010; Hüllermeier and Waegeman, 2021) and out-of-distribution detection (Yang et al., 2024; Gruber et al., 2025) could be used on the modeling side to reject predictions where the estimated uncertainty is too high for the predictions to be useful in practice.
Finally, as little is known about the composition of uncertainty in conflict modeling, more research is also needed to understand to what extent this is due to data issues and to what extent violence is the consequence of inherently stochastic processes. Much of this question relates back to conflict datasets and the challenges of recording events globally as discussed above. For example, fine-grained local datasets may report many instances of violence not contained in the larger data collection efforts needed for forecasting systems (Bazzi et al., 2022). In combination with differences in inclusion criteria and coding procedures (Raleigh et al., 2023; Öberg and Yilmaz, 2025), this leads to a large number of non-overlapping events among conflict event datasets based on very similar sources (Donnay et al., 2019), and makes it hard to establish a ground truth for further examination. Addressing this issue would go a long way towards understanding and quantifying the individual components of uncertainty in conflict forecasting, and thus enable a more targeted approach to its reduction.
6 Replication Code
The replication code for our modeling pipeline and the visualizations is available on Github at https://github.com/ccew-unibw/uncertaintrees.
7 Acknowledgements
This paper is based on a contribution to the VIEWS Prediction Challenge 2023/2024. Financial support for the Prediction Challenge was provided by the German Ministry for Foreign Affairs. For more information on the Prediction Challenge, see Hegre et al. (2025) and https://viewsforecasting.org/research/prediction-challenge-2023.
The Center for Crisis Early Warning is funded by the German Federal Ministry of Defense and the German Federal Foreign Office. The views and opinions expressed in this article are those of the authors and do not necessarily reflect the official policy or position of any agency of the German government.